OpenStudy (anonymous):

ln(2x-1)=0 simplify and solve

OpenStudy (anonymous):

x=1/2

OpenStudy (anonymous):

why tho

OpenStudy (anonymous):

u add it first then subtract

OpenStudy (amoodarya):

no x=1/2 is not define able in this log

OpenStudy (amoodarya):

ln(2x-1)=0 ln(2x-1)=ln(1) so 2x-1 =1 2x=2 x=1

OpenStudy (amoodarya):

\[\log_{b}x \rightarrow\\x>0\\b>0\\b \neq1 \]